### v1

@@ 1 Introduction
$\rho_{n}(\tau_{n}-1)=(1-\alpha)(\tau^{2}_{n}-n\tau_{n}),$
hence (14) can be written simply as
@@
Claim 5: If the equality in (14) holds, then $Tr_{min}=n-1$ and $Tr_{max}=Tr_{min}+\rho_{n}=n-1+\rho_{n}.$
Proof: If the equality in (14) holds, then the above equations hold. Furthermore, we get that
$\frac{x_{u}}{x_{v}}=\frac{x_{max}}{x_{min}}=\frac{1}{1-\tau_{n}}.$ (16)
From (13), we have $x_{w}=x_{max}$ for $w\in V(G_{0})-\{v\}$. Since $x_{v}=x_{min}$, from (12), we have
@@
Then we will proof that $Tr_{w}=Tr_{max}$ for ${w\in V(G_{0})-\{v\}}$. If it does not true, we have $\sum\limits_{w\in V(G_{0})-\{v\}}(Tr_{max}-Tr_{w})x_{max}\geq x_{max}$. From Claim 3, we get $(Tr_{max}-Tr_{v})x_{v}\geq x_{v}=x_{min}.$ Since $G_{0}$ is non-transmission regular, then we have $x_{max}>x_{min}$. Hence
@@
Hence
$\mu_{\alpha}x_{v}$ $=\alpha Tr_{v}x_{v}+(1-\alpha)\sum\limits_{w\in V(G_{0})}d(v,w)x_{w}$
$=\alpha Tr_{min}x_{v}+(1-\alpha)Tr_{min}x_{max}.$
From (16), we get
$\mu_{\alpha}=\alpha Tr_{min}+(1-\alpha)Tr_{min}\frac{x_{u}}{x_{v}}=\alpha Tr_{min}+(1-\alpha)Tr_{min}\frac{1}{1-\tau_{n}}.$ (18)
From (14), (17) and (18), we get
@@
which implies $Tr_{min}=\frac{1}{1-\alpha}\frac{n-1}{n-\tau_{n}}\frac{1-\tau_{n}}{\tau_{n}}\rho_{n}.$ Since (5), we get $\rho_{n}=\frac{(1-\alpha)\tau^{2}_{n}-(1-\alpha)n\tau_{n}}{\tau_{n}-1}.$ Then
$\begin{cases}Tr_{min}=\frac{1}{1-\alpha}\frac{n-1}{n-\tau_{n}}\frac{1-\tau_{n}}{\tau_{n}}\rho_{n}\\ \rho_{n}=\frac{(1-\alpha)\tau^{2}_{n}-(1-\alpha)n\tau_{n}}{\tau_{n}-1}\end{cases},$
by a simple calculation, we get $Tr_{min}=n-1$ and $Tr_{max}=Tr_{min}+\rho_{n}=n-1+\rho_{n}$, completing the proof of this claim.
Hence, we will characterize the necessary and sufficient conditions for the equality in (14).
On the one hand, from Claim 5, we get $d(v)=n-1$ and $Tr_{min}=Tr_{v}=n-1$, which implies $dim(G)=2$. For any $y\in V(G_{0})$,
@@
Hence, for $z\in V(G_{0})-\{v\}$, we have $d(z)=2(n-1)-Tr_{z}=2n-2-(n-1+\rho_{n})=n-1-\rho_{n}$, which implies that if $n$ is odd, $d(z)=n-2$; if n is even, $d(z)=n-3.$ Hence, if $n$ is odd, then $G_{0}\cong K_{1,2,\cdots,2}$; if $n$ is even, then $G_{0}$ is isomorphic to any $(n-4)$-$DVDR$ graph.
On the other hand, under the following conditions:
(1) If $n$ is odd, $\alpha\in[0,1)$ and $G\cong K_{1,2,\cdots,2}.$
(2) If $n$ is even, $\alpha\in[0,1)$ and $G_{0}$ is isomorphic to any $(n-4)$-$DVDR$ graph.
It is easy to see that
$Tr_{max}(G)-\mu_{\alpha}(G)=\frac{\rho_{n}(1-\tau_{n})}{n-\tau_{n}}=(1-\alpha)\tau_{n},$
where $\tau_{n}=\tau_{n}(1)$ if $n$ is odd and $\tau_{n}=\tau_{n}(2)$ if $n$ is even. Therefore, the equation clearly holds.
Since $G_{0}$ is a graph which attains the minimum of $Tr_{max}(G)-\mu_{\alpha}(G)$ among all connected non-transmission regular graph $G$ of order $n$. Hence, for any connected non-transmission regular graph $G$, we have:
$Tr_{max}(G)-\mu_{\alpha}(G)\geq Tr_{max}(G_{0})-\mu_{\alpha}(G_{0})\geq(1-\alpha)\tau_{n}.$
@@
where $\tau_{n}(1)=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2(1-\alpha)}.$ Equality holds if and only if $G\cong K_{1,2,\cdots,2}.$ (2) If $n$ is even and $\alpha\in[0,1)$, then
@@
where $\tau_{n}(2)=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2(1-\alpha)}.$ Equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph. ∎
In the following, let $\alpha=0$, then $\tau_{n}(1)=\frac{n+1-\sqrt{(n+1)^{2}-4}}{2}$ and $\tau_{n}(2)=\frac{n+2-\sqrt{(n+2)^{2}-8}}{2}.$ Hence, we can get the same conclusion as Theorem 1.1 in [15].
Corollary 2.7.
([15]) Let $G$ be a connected non-transmission regular graph of order $n$. Let $\mu(G)$ be the distance spectral radius of $G$, then (1) If $n$ is odd, then
$Tr_{max}-\mu(G)\geq\frac{n+1-\sqrt{(n+1)^{2}-4}}{2}.$ (19)
Equality holds if and only if $G\cong K_{1,2,\cdots,2}.$ (2) If $n$ is even, then
$Tr_{max}-\mu(G)\geq\frac{n+2-\sqrt{(n+2)^{2}-8}}{2}.$ (20)
Equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph.
In the following corollary, let $\alpha=\frac{1}{2}$, then $\tau_{n}(1)=\frac{n+2-\sqrt{n^{2}+4n-4}}{2}$ and $\tau_{n}(2)=\frac{n+4-\sqrt{n^{2}+8n}}{2}.$ Hence, we can get the same conclusion as Theorem 1.1 in [11].
Corollary 2.8.
([15]) Let $G$ be a connected non-transmission regular graph of order $n$. Let $\mu_{Q}(G)$ be the distance signless Laplacian spectral radius of $G$, then (1) If $n$ is odd, then
$2Tr_{max}-\mu_{Q}(G)\geq\frac{n+2-\sqrt{n^{2}+4n-4}}{2}.$ (21)
Equality holds if and only if $G\cong K_{1,2,\cdots,2}.$ (2) If $n$ is even, then
$2Tr_{max}-\mu_{Q}(G)\geq\frac{n+4-\sqrt{n^{2}+8n}}{2}.$ (22)
Equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph.

### v2

@@ 2 Preliminaries
$\rho_{n}(\tau_{n}-1)=(1-\alpha)(\tau^{2}_{n}-n\tau_{n}).$
Furthermore, (15) can be written simply as
@@
By (7), noting that $Tr_{max}-\mu_{\alpha}\leq(1-\alpha)\tau_{n}$.
Hence
$Tr_{max}-\mu_{\alpha}=(1-\alpha)\tau_{n},$
and all inequalities above must be equalities.
Claim 5. $Tr_{min}=n-1$ and $Tr_{max}=Tr_{min}+\rho_{n}=n-1+\rho_{n}.$
Proof. According to the proof of Claim 4, we know that all the inequalities in Claim 4 must be equalities.
From the equality in (14), it can be concluded that $x_{w}=x_{max}$ for $w\in V(G_{0})-\{v\}$. Since $x_{v}=x_{min}$, from the equality in (13), we have
@@
Next, we will proof that $Tr_{w}=Tr_{max}$ for $w\in V(G_{0})-\{v\}$}.
If it does not true, we have $\sum\limits_{w\in V(G_{0})-\{v\}}(Tr_{max}-Tr_{w})x_{max}\geq x_{max}$. From Claim 3, we know $(Tr_{max}-Tr_{v})x_{v}\geq x_{v}=x_{min}.$ Since $G_{0}$ is non-transmission regular, then $x_{max}>x_{min}$. Hence
@@
Furthermore
$\mu_{\alpha}{x_{min}}=\mu_{\alpha}x_{v}$ $=\alpha Tr_{v}x_{v}+(1-\alpha)\sum\limits_{w\in V(G_{0})}d(v,w)x_{w}$
$=\alpha Tr_{min}{x_{min}}+(1-\alpha)Tr_{min}x_{max}.$
From Claim 2, $\frac{x_{max}}{x_{min}}=\frac{1}{1-\tau_{n}}$, then
$\mu_{\alpha}=\alpha Tr_{min}+(1-\alpha)Tr_{min}\frac{x_{max}}{x_{min}}=\alpha Tr_{min}+(1-\alpha)Tr_{min}\frac{1}{1-\tau_{n}}.$ (18)
From the equality in (15), (17) and (18), it can be concluded that
@@
which implies $Tr_{min}=\frac{1}{1-\alpha}\frac{n-1}{n-\tau_{n}}\frac{1-\tau_{n}}{\tau_{n}}\rho_{n}$.
Recall that $\tau_{n}$ satisfies the equation (6), we get $\rho_{n}=\frac{(1-\alpha)(\tau^{2}_{n}-n\tau_{n})}{\tau_{n}-1}.$ Then
$\begin{cases}Tr_{min}=\frac{1}{1-\alpha}\frac{n-1}{n-\tau_{n}}\frac{1-\tau_{n}}{\tau_{n}}\rho_{n}\\ \rho_{n}=\frac{(1-\alpha)(\tau^{2}_{n}-n\tau_{n})}{\tau_{n}-1}\end{cases}.$
By a simple calculation, we get $Tr_{min}=n-1$.
Furthermore, by (17), we have $Tr_{max}=Tr_{min}+\rho_{n}=n-1+\rho_{n}$, which completes the proof of this claim.
According to the proof process mentioned above, we continue to prove Theorem 1.1.
From Claim 5, we get $Tr_{min}=Tr_{v}=n-1$, which implies d(v)=n1d(v)=n-1 and $diam(G_{0})=2$. For any $y\in V(G_{0})$,
@@
Hence, for $z\in V(G_{0})-\{v\}$, we have $d(z)=2(n-1)-Tr_{z}=2n-2-(n-1+\rho_{n})=n-1-\rho_{n}$, which implies that if $n$ is odd, d(z)=n2d(z)=n-2; if $n$ is even, d(z)=n3.d(z)=n-3. Hence, if $n$ is odd, then $G_{0}\cong K_{1,2,\cdots,2}$; if $n$ is even, then $G_{0}$ is isomorphic to any (n4)(n-4)-DVDRDVDR graph.
Note that $G_{0}$ is a graph which attains the minimum $Tr_{max}(G)-\mu_{\alpha}(G)$ among all connected non-transmission regular graph $G$ of order $n$. Hence, for any connected non-transmission regular graph $G$ of order $n$, it can be concluded that:
$Tr_{max}(G)-\mu_{\alpha}(G)\geq Tr_{max}(G_{0})-\mu_{\alpha}(G_{0})=(1-\alpha)\tau_{n}.$
@@
where $\tau_{n}(1)=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2(1-\alpha)}$. Equality holds if and only if $G\cong K_{1,2,\cdots,2}$.
(2) If $n$ is even and $\alpha\in[0,1)$, then
@@
where $\tau_{n}(2)=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2(1-\alpha)}$. Equality holds if and only if $G$ is isomorphic to any (n4)(n-4)-DVDRDVDR graph. ∎
Statement
This article has undergone further revisions and enhancements, all of which were contributed by Ligong Wang. In recognition of his contributions, Ligong Wang is now acknowledged as a new co-author. The manuscript has been updated to its current version to incorporate these changes. We affirm that all authors have reviewed and approved this update.